Notes

Sort by:
  • Reference Designator
  • First Name
  • Created
  • Modified  
  • Metadata Start Date End Date Comment
    GP05MOAS-GL276-01-FLORDM000
    Deployments 1 and 2: Data streams for this instrument will not match what we have in the database!! A FLORT was deployed (not a FLORD). The data had to be ingested using the FLORTM driver/parser, meaning the data streams are flort_sample.

    By Lori Garzio, on 8/20/18
    Redmine Issue #13507

    CE01ISSM-MFD37-03-DOSTAD000
    Deployment: 8
    Asset ID: CGINS-DOSTAD-00483
    Status: Open Issue
    10/12/17 4/15/18 A recent routine control check of the O2 reference sensors, used for multipoint calibration of oxygen Optodes by the manufacturer, revealed a deviation with respect to accuracy. Regrettably this led us to conclude that all new and newly repaired multipoint calibrated oxygen optodes delivered from Aanderaa Norway between January 4th 2017 and October 20th 2017 are outside of specification. 3 Endurance DOSTAs are affected, two have been deployed. 478 - in inventory, will be sent back to...

    By Mike Smith, on 1/18/18
    Redmine Issue #13080

    CE06ISSM-MFD37-03-DOSTAD000
    Deployment: 7
    Asset ID: CGINS-DOSTAD-00479
    Status: Open Issue
    10/5/17 4/15/18 A recent routine control check of the O2 reference sensors, used for multipoint calibration of oxygen Optodes by the manufacturer, revealed a deviation with respect to accuracy. Regrettably this led us to conclude that all new and newly repaired multipoint calibrated oxygen optodes delivered from Aanderaa Norway between January 4th 2017 and October 20th 2017 are outside of specification. 3 Endurance DOSTAs are affected, two have been deployed. 478 - in inventory, will be sent back to...

    By Mike Smith, on 1/18/18
    Redmine Issue #13080

    CE07SHSM-MFD37-01-OPTAAD000
    Status: Not Operational
    10/2/17 4/2/18 MFN OPTAA not deployed OPTAA on MFN not deployed because instruments were not serviced in time by the vendor. We prioritized the NSIF OPTAAs over the MFN ones because of greater phytoplankton concentrations at the NSIF depth

    By Mike Smith, on 12/6/17
    Redmine Issue #12982

    CE01ISSM-MFD37-01-OPTAAD000
    Status: Not Operational
    10/2/17 4/2/18 MFN OPTAA not deployed OPTAA on MFN not deployed because instruments were not serviced in time by the vendor. We prioritized the NSIF OPTAAs over the MFN ones because of greater phytoplankton concentrations at the NSIF depth

    By Mike Smith, on 12/6/17
    Redmine Issue #12982

    CE06ISSM-MFD37-01-OPTAAD000
    Status: Not Operational
    10/2/17 4/2/18 MFN OPTAA not deployed OPTAA on MFN not deployed because instruments were not serviced in time by the vendor. We prioritized the NSIF OPTAAs over the MFN ones because of greater phytoplankton concentrations at the NSIF depth

    By Mike Smith, on 12/6/17
    Redmine Issue #12982

    CE04OSSM-SBD11-06-METBKA000
    Deployment: 5
    Asset ID: CGINS-METLGR-00030
    Status: Open Issue
    11/14/17 4/2/18 CE04OSSM METBK wind sensor stopped working The sensor stopped outputting data on Nov-14. This changes the data stream from the METBK, so no METBK sensor data are being read properly by CI from this mooring. The other METBK sensors continue to output reasonable data.

    By Mike Smith, on 12/6/17
    Redmine Issue #12961

    CE02SHSM-RID26-04-VELPTA000
    Deployment: 6
    Asset ID: CGINS-VELPTA-11775
    Status: Open Issue
    11/20/17 4/2/18 VELPT on CE02SHSM NSIF tilts too much during storms NSIF tilted over to ~30 degrees according to its VELPT. According to VELPT manufacturer, data are suspect when this instrument is tilted more than 20 degrees. A quality flag should be set when this happens. Pressure sensor indicates the NSIF shallowed during this time. Our assumption is that the mooring is stretched taut in one direction during storms. One should also check MOPAK and VELPT-buoy data to see if the buoy tilts excessively...

    By Mike Smith, on 12/6/17
    Redmine Issue #12960

    CE01ISSM-RID16-01-OPTAAD000
    Deployment: 8
    Asset ID: CGINS-OPTAAD-00149
    Status: Failed
    9/30/17 12/6/17 CE01ISSM NSIF OPTAA stopped working CE01ISSM NSIF OPTAA port stopped recording full data records in the last few days of October. This problem got worse over the week until no data were coming through. On 11/6 the OPTAA port was turned off.

    By Mike Smith, on 12/6/17
    Redmine Issue #12959

    CE01ISSM
    Deployment: 8
    Asset ID: CGMCE-01ISSM-00008
    Status: Failed
    10/24/17 4/2/18 DSL on CE01ISSM stopped working We experienced poor communications via the DSL to the MFN after hooking the system up on deck (10/12). While we were able to adjust the communication rate to correct for this, subsequently (10/18) there have been times when the surface buoy did not connect to the MFN via the DSL. This is how data from the MFN are telemetered to the buoy. All sensors on the MFN appear to be working based on power consumption data, but science data transfers from the MFN...

    By Mike Smith, on 12/6/17
    Redmine Issue #12958

    CE01ISSM-MFD35
    Status: Failed
    10/24/17 4/2/18 DSL on CE01ISSM stopped working We experienced poor communications via the DSL to the MFN after hooking the system up on deck (10/12). While we were able to adjust the communication rate to correct for this, subsequently (10/18) there have been times when the surface buoy did not connect to the MFN via the DSL. This is how data from the MFN are telemetered to the buoy. All sensors on the MFN appear to be working based on power consumption data, but science data transfers from the MFN...

    By Mike Smith, on 12/6/17
    Redmine Issue #12958

    CE01ISSM-MFD37-06-CAMDSA000
    Deployment: 8
    Asset ID: CGINS-CAMDSA-00110
    Status: Failed
    10/14/17 4/2/18 CE01ISSM CAMDS stopped working The CAMDS experienced some kind of failure wherein the sampling schedule suddenly changed from every 4 hours to approximately every 5 minutes. The camera stopped taking pictures 10/14 after either blowing through its batteries or disk space. This problem has not been seen before.

    By Mike Smith, on 12/6/17
    Redmine Issue #12957

    CE07SHSM-SBD11-06-METBKA000
    Deployment: 6
    Asset ID: CGINS-METLGR-00029
    Status: Failed
    10/11/17 4/2/18 CE07SHSM METBK wind sensor stopped working Wind sensor stopped working. This means the data stream from the METBK is changed, so no data from any of the METBK sensors are being read properly by CI. However, all of these other METBK are actually working and telemetering reasonable data. METBK parsers cannot read the METBK data from this instrument now because the DCL changed the format of the file upon wind sensor failure.

    By Mike Smith, on 12/6/17
    Redmine Issue #12963

    CE07SHSM-RID27-01-OPTAAD000
    Deployment: 6
    Asset ID: CGINS-OPTAAD-00171
    Status: Failed
    10/24/17 4/2/18 CE07SHSM NSIF OPTAA stopped working Deployed Oct-3. Oct-24 it was observed to no longer be delivering full 3 minute bursts of data to its DCL. It got progressively worse, so on Nov-2 it was turned off.

    By Mike Smith, on 12/6/17
    Redmine Issue #12964

    CE07SHSM
    Deployment: 6
    Asset ID: CGMCE-07SHSM-00006
    Status: Open Issue
    11/2/17 4/2/18 CE07SHSM no longer able to recharge batteries (PSC) The mooring stopped receiving power from photo voltaic panel #1 of 4 on Nov-2. By Nov-20, the mooring stopped receiving power from two more solar panels (PV2 & PV4) and one of two wind turbines (WT2). This indicates the Power System Controller (PSC) they connect to is failing. Mooring components will be shut down as the mooring loses the ability to recharge its batteries.

    By Mike Smith, on 12/6/17
    Redmine Issue #12965

    CE09OSSM-MFD35-04-ADCPSJ000
    Deployment: 6
    Asset ID: CGINS-ADCPSJ-19061
    Status: Failed
    10/9/17 4/2/18 On 10/9/2017, 4 days after deployment, it sent the following error message: 2017/10/09 02:32:36.797 [BREAK Wakeup A] 2017/10/09 02:32:36.839 WorkHorse Broadband ADCP Version 50.40 2017/10/09 02:32:36.880 Teledyne RD Instruments (c) 1996-2010 2017/10/09 02:32:36.903 All Rights Reserved. 2017/10/09 02:32:36.962 ERR: No frequency information available! 2017/10/09 02:32:37.014 Please use &Q command to enter system frequency. 2017/10/09 02:32:37.062 Setting System Frequency To Default 300...

    By Mike Smith, on 12/6/17
    Redmine Issue #12978

    CE09OSSM-RID26-01-ADCPTC000
    Deployment: 6
    Asset ID: CGINS-ADCPTC-22809
    Status: Open Issue
    10/6/17 4/2/18 ADCPT started receiving/generating random break commands right after deployment. Interventions at the time were unable to resolve, but we believe the system was collecting data. We reset it and disconnected it from the DCL. Hopefully, it will run autonomously. We presume the problem was caused by the DCL sending erroneous commands, but it could be something else. We’ll see after recovery. Telemetered data no longer flowing to the system. Data should be available upon recovery.

    By Mike Smith, on 12/6/17
    Redmine Issue #12979

    CE09OSSM
    Deployment: 6
    Asset ID: CGMCE-09OSSM-00006
    Status: Not Operational
    10/11/17 4/2/18 10/11 lost wind turbine 2 (out of 2) 10/24 batteries not recharging sufficiently. To conserve power, the MFN was turned off and the NSIF was reduced to 1 hour communications. The OPTAA is the only instrument on the MFN without internal batteries. 11/8 restored power to entire mooring after a period of strong wind (utilizing WT1)

    By Mike Smith, on 12/6/17
    Redmine Issue #12980

    CE06ISSM
    Deployment: 7
    Asset ID: CGMCE-06ISSM-00007
    Status: Not Operational
    10/3/17 4/2/18 CE06ISSM-MFD37-06-CAMDSA CAMDS not serviced by vendor in time for deployment

    By Mike Smith, on 12/6/17
    Redmine Issue #12981

    CE02SHSM-SBD12-08-FDCHPA000
    Deployment: 6
    Asset ID: CGINS-FDCHPA-00024
    Status: Not Operational
    10/15/17 4/2/18 - CE02SHSM deployed Oct 10. The Power System Controller (PSC) failed 9 hours later. - CE02SHSM recovered Oct 14. Buoy from CE02SHSM swapped with recently recovered CE04OSSM. CE02SHSM's refurbished instrument plate was moved to the recently recovered CE04OSSM buoy, so the hybrid buoy to be deployed had a newly refurbished METBK-CT and PCO2A. Other METBK sensors from previously deployed CE04OSSM remained, but they were all still working and their calibrations would still be valid through...

    By Mike Smith, on 12/6/17
    Redmine Issue #12984

    1 of 11